Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #5808

Period Covered:
08:00 pm June 7, 2012 - 07:59 pm June 8, 2012
(DOY 160/0000z - 160/2359z)

Flight Operations Summary

This report was generated by the HST Automated Report System @2012-06-09 17:02:00 GMT

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)




FGS GSAcq		9		9
FGS REAcq		8		8
OBAD with Maneuver	8		8

The COS FUV Special Recovery activities (proposal 12810 visits 22 - 23) successfully completed 
today without incident.  The FUV was ramped to -5293V for segment A (178 counts) and -5246V (175 
count) for B.  The QE grid was off.   No elevated counts or HV current transients were observed in 
the engineering telemetry during or after the ramp.  The STScI and MOSES2 Engineering Teams have 
examined the relevant telemetry and confirmed all voltages, currents and temperatures were as 
expected. The voltage and current monitors matched the signatures seen in SMOV.  The COS Science 
Team at STScI has examined the science data and DCE memory dumps and everything is nominal.  No HV 
current transients or unexpected elevated count rates were observed in the DCE memory dumps. The 
science data consisted of 2 sets of dark and wavecal exposures. The exposures were collected 
immediately after ramping to HVNom (178/175).  Comparisons with SMOV data taken indicates 
comparable behavior when the aging of the detector, QE grid off,  and temporal fluctuations are 
considered. The Engineering and Science Teams have given a "Go" to proceed with visit 24 and 25. 
Visits 24 and 25 are scheduled for Monday June 11 (DOY 163).  The COS event flag 3 was cleared 
@160/17:27:28 via OR 19268  

Program Principal Investigator Program Title
Science Observations Scheduled
12210 Adam Bolton, University of Utah SLACS for the Masses: Extending Strong Lensing to Lower Masses and Smaller Radii
12473 David Sing, University of Exeter An Optical Transmission Spectral Survey of hot-Jupiter Exoplanetary Atmospheres
12488 Mattia Negrello, Open University SNAPshot observations of gravitational lens systems discovered via wide-field Herschel imaging
12493 Ian McGreer, University of Arizona A Candidate Lensed Quasar at z=6.25
12517 Francesco Ferraro, Universita di Bologna COSMIC-LAB: Hunting for optical companions to binary MSPs in Globular Clusters
12521 Xin Liu, University of California - Los Angeles The Frequency and Demographics of Dual Active Galactic Nuclei
12533 Crystal Martin, University of California - Santa Barbara Escape of Lyman-Alpha Photons from Dusty Starbursts
12659 Joaquin Vieira, California Institute of Technology Strongly Lensed Dusty Star Forming Galaxies: Probing the Physics of Massive Galaxy Formation
Calibration Observations Scheduled
12689 Tiffany Borders, Space Telescope Science Institute WFC3 UVIS CCD Daily Monitor
12690 Cheryl Pavlovsky, Space Telescope Science Institute UVIS Gain Stability
12732 David Golimowski, Space Telescope Science Institute CCD Hot Pixel Annealing
12742 Justin Ely, Space Telescope Science Institute CCD Dark Monitor Part 2
12744 Justin Ely, Space Telescope Science Institute CCD Bias and Read Noise Monitor-Part 2
12782 David Golimowski, Space Telescope Science Institute CCD Daily Monitor {Part 2}
Other Observations Scheduled
12810 Thomas Wheeler, Space Telescope Science Institute COS/FUV Special Recovery